OpenStudy (anonymous):

what value of x satisfies the equation 2/3x + 1 = 9?

OpenStudy (♪chibiterasu):

Please don't give out straight answers.

OpenStudy (anonymous):

show her how to do it...don't just give answers

OpenStudy (anonymous):

how u got tht

OpenStudy (anonymous):

First, simplify the equation by subtracting 1 by both sides. 2/3x = 8 Now divide both sides by 2/3 to isolate x x = 8/1 / 2/3 To divide fractions multiply the first fraction by the reciprocal of the second fraction 8/1 * 3/2 8*3 = 24 1*2 = 2 24/2 This can be simplified by dividing both numerator and denominator by 2, giving you 12/1, which then simplifies to 12

OpenStudy (♪chibiterasu):

First we just subtract one from both sides in order to cancel out: 2/3x + 1 - 1 = 9 - 1 2/3x = 8 Now we have to divide 2/3 from both sides to cancel out as well: 2/3x/(2/3) = 8/(2/3) But first let's divide 2/3: 0.67x/0/67 = 8/0.67 x = 11.94 which is rounded to 12.

OpenStudy (anonymous):

ok got it !!! help me with another?

OpenStudy (♪chibiterasu):

Lol yeah that's another way to do it.

OpenStudy (austinl):

We have, \(\displaystyle \frac{2}{3}x+1=9\) First we want to get the terms with variables on one side and constants on the other, correct? So we would subtract 1 from each side. \(\displaystyle \frac{2}{3}x=8\) To get rid of the fraction we need to multiply by it's reciprocal \(\frac{3}{2}\) \(\displaystyle\frac{3}{2}\cdot\frac{2}{3}x=8\cdot\frac{3}{2}\) \(x=\dfrac{8\times3}{2}=\dfrac{24}{3}\) When simplified that is equal to 12. THAT is how you get to that answer.

OpenStudy (anonymous):

if 3b + 12 = 5(b-4), what is b?

OpenStudy (austinl):

For this one, you would need to distribute out the right side. Do you think you could show me what that looks like?

OpenStudy (anonymous):

would it be .... b+8=5?

OpenStudy (austinl):

No to distribute you need to do something like this. \(a(b-c)=ab-ac\) Make sense?

undeadknight26 (undeadknight26):

Austin is that ur mustang?

OpenStudy (austinl):

Two things, 1) Not really the place for that. You shouldn't just interrupt someone's question for something off topic. 2) Yes it is.

OpenStudy (♪chibiterasu):

Let's move on :I

OpenStudy (anonymous):

idk how to do tht

OpenStudy (austinl):

We have, \(5(b-4)\) This would translate to, \(5(b-4)\Rightarrow5b-(5\cdot4)=5b-20\) So all told we now have, \( 3b + 12 =5b-20\) Do you think you could get terms with variables on one side and constants on the other?

OpenStudy (anonymous):

lol no im lost , wht you done i was doin completely different i did 3(b+4_=3b+8

OpenStudy (austinl):

Where did you get these numbers from??

OpenStudy (anonymous):

24

OpenStudy (anonymous):

noo 8

OpenStudy (anonymous):

3b + 12 = 5(b - 4) Distribute the 5 on the right side to simplify it 3b + 12 = 5b - 5(4) 3b + 12 = 5b - 20

OpenStudy (anonymous):

Can you go from there?

OpenStudy (anonymous):

3b + 12 = 5(b - 4) 3b + 12 = 5 * b + 5*-4 3b + 12 = 5b + (-20) 3b + 12 = 5b - 20 you understand so far ?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

No so if you have 3b + 12 = 5b - 20 You'd want to get like terms on one side, so b on the left and and natural numbers on the right, it will be easier to do it this way To do this you want to -12 from both sides of the equation like this: 3b +12 = 5b - 20 -12 -12 3b = 5b - 32 When you subtracted 12 from both sides, you want to keep the - in front of 20, so you are subtracting 12 from -20, which gives you -32, then in conclusion, when placed back in the equation simplifies to 3b = 5b - 32 Try to get b on one side by itself now

OpenStudy (anonymous):

do i add 3 and 5?

OpenStudy (anonymous):

because 5b is a positive number imagine a + in front of it, you want to always reverse the operation when moving to the other side, so you want to subtract 5b from 3b

OpenStudy (anonymous):

3b-5b = -2b Now you have -2b = -32 we want to find b by itself so divide both sides by -2 -2b / -2 = b -32 / -2 = +16 b = 16 to check this replace b with 16 in the equation 3*16 + 12 = 5*16 - 20 48 + 12 = 80 - 20 60 = 60 the equation is true so b = 16

OpenStudy (anonymous):

i get it ... what bout .. what value of g make 3g+9=18 a true statementt

OpenStudy (anonymous):

first you want to get 3g on one side and everything else on the other side. 3g + 9 = 18 (subtract 9 from both sides) 3g + 9 - 9 = 18 - 9 (what you do one one side of the equal sign, you have to do on the other side) now we have : 3g = 9 are you still with me ?

OpenStudy (anonymous):

hello ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

3g = 9 now we will divide both sides by 3 to get g by itself 3g/3 = 9/3 g = 3 the value of g being 3 makes it a true statement

OpenStudy (anonymous):

thanks ! what about what is the solution to the equation below? 5/4x - 5/2 = 2